OpenStudy (anonymous):

whats p=2(l+w) for l

OpenStudy (texaschic101):

p = 2(l + w) p = 2l + 2w --- subtract 2w from both sides p - 2w = 2l -- divide both sides by 2 (p - 2w) / 2 = l or p/2 - w = l